OpenStudy (anonymous):

Use an inequality to solve the problem. Be sure to show the inequality and all of your work. One side of a rectangle is 4 times the other, and the perimeter is not to exceed 150 inches. Find the possible values for x, the length of the shorter side.

OpenStudy (anonymous):

2(4x+x)=15 8x+2x=150 150\10=15 x=15

OpenStudy (anonymous):

This an inequality not an equation so ... side 1 is given... x Side 2 is given... 4x Perimeter therefore is given ... 2( x + 4x) = 10x Inequality therefore is given ... 10x<150 divide both sides by 10 ... x<15 Voila!